OpenStudy (anonymous):

Write the equation of the line below in slope intercept form. Points: (-5,0) (-3,-1) (-2,0) and (3,-3) I don't understand how to do this?

hero (hero):

Have you plot the points on graph paper yet?

OpenStudy (anonymous):

Yes!

hero (hero):

You have four points that are not in a straight line. You only need two points to figure out the line. So which two points go together?

OpenStudy (anonymous):

Wait, they are in a straight line! It slants like \

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

@JakeV8 Help please?

hero (hero):

They're not in a straight line. Maybe you posted the points incorrectly. Double check each point. Make sure all of the signs are correct for each coordinate.

OpenStudy (anonymous):

The first one should be (-6,0) !! My bad

hero (hero):

Make sure the rest of them are correct as well.

hero (hero):

So far, they are still not lining up

OpenStudy (anonymous):

They are! I just checked!(:

hero (hero):

Nope, they are not.

OpenStudy (anonymous):

I have a picture but it won't let me post it!:/

hero (hero):

I assume your picture plots the points but leaves it up to you to label them, right?

OpenStudy (anonymous):

 That's what it looks like! and no, it's already labeled!m Let me check the points one more time

OpenStudy (anonymous):

(-6,0) (-3,-1) (0,-2) (3,-3) Does that make more sense?

hero (hero):

Yes, finally

hero (hero):

Well, from the graph, you can easily figure out the slope and y-intercept

hero (hero):

In fact, you should already know what the y-intercept is

OpenStudy (anonymous):

The y-intercept would be (0,-2) correct?

hero (hero):

Yes, correct or just simply b = -2

hero (hero):

What is the slope?

OpenStudy (anonymous):

Slope is up and over right? or does that really matter?

hero (hero):

slope = rise over run The first thing you should do is draw a line through the points and ask yourself is the line rising or falling. If it is rising, then the slope is positive. If it is falling, then the slope is negative. Let me know.

OpenStudy (anonymous):

It is falling so it would be -1/3 right? It rises one and runs three

hero (hero):

So now that you have the slope and y-intercept, insert m = -1/3 and b = -2 into the formula y = mx + b.

OpenStudy (anonymous):

y=-1/3+(-2) ? or would it turn negative like y=-1/3-2 That's the part that confuses me.

hero (hero):

y = -1/3(x) - 2

hero (hero):

Or you can write it as: y = -x/3 - 2

OpenStudy (anonymous):

Okay! Thank you SO SO much for working with me here!